Three consecutive integers add up to 51. What are these integers? 

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Let the three consecutive integers be x, x+1 and x+2.
According to the question, x+(x+1)+(x+2)=51
⇒3x+3=51
⇒3x=51-3
⇒3x=48
⇒x=48/3
⇒x=16
Thus, the integers are x=16
x+1=17 x+2=18

Thanks